# HLAT: High-quality Large Language Model Pre-trained on AWS Trainium

### v1

@@ 5.3. Truthfulness and Bias
We report the model’s truthfulness and bias using TruthfulQA (Lin et al., 2022) and CrowS-pairs (Nangia et al., 2020). TruthfulQA presents a collection of meticulously crafted questions spanning diverse domains such as health, law, finance, and even politics. These queries deliberately target areas where human intuition and personal biases can lead to incorrect responses, and measure an LLM’s resistance to misinformed or erroneous knowledge. CrowS-Pairs is a benchmark designed to probe LLMs for social biases across nine categories, including gender, religion, race/color, sexual orientation, age, nationality, disability, physical appearance and socioeconomic status. Each example is composed of a stereotype and an anti-stereotype.
@@
We present the results in Table 4 with 0 shot inference. For TruthfulQA, we measure the multiple-choice score, and higher score shows better truthfulness. For CrowS-Pairs, it measures the percentage of models choosing answers of stereotypes, so lower scores indicates smaller bias. Overall, HLAT  performs similar as other opensource models.
5.4. Efficiency and Scalability
We describe the training efficiency in terms of Cost per 4-million tokens (CPT) and scalability reported in (N.A., [n. d.]). For comparison, the GPU baseline is established using p4d.24xlarge instances and NeMo 23.08 (Harper et al., [n. d.]) (available inside NeMo docker container with tag 23.08) software stack. Figure 4 plots the normalized CPT of training on Trainium  and scaling. For reference, we include both a 7B model and a 70B model (similar architecture as LLaMA-70B (Touvron et al., 2023b)). The Trainium  CPT is normalized, such that the CPT of the GPU baseline is 100%. On 64 nodes, training HLAT  7B model costs about 54% of the GPU baseline.
@@ 5.4. Efficiency and Scalability
Figure 4. Normalized cost per 4 million tokens (CPT) for 7B and 70B models on AWS Trainium with various number of nodes. CPT of GPU baseline is normalized to 100%. 70B models ran into out-of-memory on 4 nodes.
5.5. Model Limitation
@@
We note some limitations of HLAT in this section. Similar as other LLMs, HLAT suffers a set of limitations such as hallucinations, potential non-factual generations, biases, and toxicity (Zhang et al., 2023). For example, although comparable with other open-source pre-trained models, the bias of HLAT is still relative high on some subjects such as sexual orientation, physical appearance, religion, and socioeconomic (see Table 4). This is partially due to the usage of publicly available datasets. More importantly, as a pre-trained model HLAT  has not gone through a supervised finetuning and human preference alignment. Those fine-tuning methods have been shown to be able to alleviate some limitations of pre-trained LLMs (Touvron et al., 2023b). Another limitation is our training is stopped after 1.8 trillion tokens. As is suggested by Figure 3, HLAT may be able to further improve on certain tasks, such as math and world knowledge, with more training tokens.
6. Best Practices & Future Directions
@@
Parallelism: NDTL supports TP up to 32 degrees and pipeline parallelism. For a 7B model, we found that the combination of TP=8 and PP=1 provides the highest training throughput. However, for models of other sizes and architectures, the optimal parallelism configuration could vary. To achieve the highest training throughput, parallelism configuration needs to be jointly optimized with choice of activation checkpointing method, gradient accumulation steps, and training precision, to avoid out-of-memory on Trainium.
Training Precision: NDTL supports various training precision settings, including full precision (FP32), AMP BF16 mixed precision, BF16 with and without SR, etc. Full precision training is often memory-wise infeasible for multi-billion LLMs. We compared three training strategies for HLAT: pure BF16, BF16 with SR, and mixed precision training. Empirically, we found that training loss of pure BF16 diverges. BF16 with SR and mixed precision get similar training loss. We finally chose BF16 with SR for higher throughput. However, for models of other sizes and architecture, BF16 with SR may not guarantee the same convergence as mixed precision. Usually, the divergence can be observed in first few thousands of steps.
@@
Weight decay: We applied weight decay to all layers. Empirically, weight decay is not applied to normalization and bias layers (Devlin et al., 2019). In our experiment, we did not found much performance-wise difference of those two methods.
Dataloader: Dataloader is another key factor in model performance. We compared our dataloader described in Section 3.3 with Nemo dataloader (Harper et al., [n. d.]) to demonstrate better performance on our dataloader. The Nemo dataloader tends to overfit due to repetition within sequence.
@@ 7. Related Work
LLM pre-training: After the Transformer architecture (Vaswani et al., 2017) was introduced, BERT (Devlin et al., 2019) was proposed to pre-train a language model on a large corpus of unlabeled data. Following the success of BERT model on various NLP tasks, many pre-trained language models are later introduced with different architectures and training methods, such as GPT-2 (Radford et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), and so on (Zhao et al., 2023b). Studies later observed significant performance improvement of language models by increasing model size and training data (Hoffmann et al., 2022). Such abilities are further demonstrated in LLMs such as GPT-3 (Brown et al., 2020), PaLM (Anil et al., 2023), LLaMA (Touvron et al., 2023a), Falcon (Almazrouei et al., 2023), etc. Pre-trained on trillions of tokens, LLMs with tens or hundreds of billions parameters show remarkable ability in generating creative text contents, as well as a variety of downstream tasks, such as question answering, summarization, machine translation, programming, etc. (Zhao et al., 2023b).
@@
AI accelerators: Most models are trained on NVIDIA GPU accelerators, such as GPT (Brown et al., 2020; OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b). Falcon-180B (Almazrouei et al., 2023) was trained on AWS SageMaker, with up to 4,096 A100 40GB GPUs using p4d instances. However, the landscape of hardware accelerators for deep learning training has blossomed in recent years, with established players like NVIDIA GPUs facing fierce competition from custom offerings like Google’s TPU and AWS Trainium. PaLM-2 (Anil et al., 2023) and OpenLLaMA (Geng and Liu, 2023) have demonstreated successful LLM pre-training on Google TPU. Recently, OLMo (Groeneveld et al., 2024) is an opensource model developed by AI2. It has two models trained on AMD and Nvidia GPUs, separately. The two models have nearly identical performance on their evaluation suite by 2T tokens. AWS Trainium is a machine learning accelerator developed for deep learning training with high performance and cost-competitiveness. Our work is the first demonstration of end-to-end multi-billion LLM pre-trained on AWS Trainium. Ultimately, the optimal choice depends on the specific needs of the training task, with further research required to fully explore the potential of each accelerator and their possible convergence in future architectures.
8. Conclusion
@@
In this paper, we pre-train HLAT, a 7 billion parameter large language model, using AWS Trainium over 1.8 trillion tokens. HLAT follows the decoder-only architecture and is trained with 64 AWS trn1.32xlarge instances. We evaluate the performance of HLAT against popular open-source baseline models including LLaMA and OpenLLaMA on a variety of popular benchmarking tasks. We find that HLAT achieves model quality on par with these baseline models. This work for the first time demonstrates that AWS Trainium with NDTL is able to successfully pre-train high-quality LLMs with high efficiency and low cost.
Appendix A Detailed Evaluation Results
In Table 5 we show the evaluation results on all individual tasks described in Section 5. The number of shots are same as those listed in Table 3.
Table 5. Evaluation results on all tasks of HLAT  (using final 1.8 trillion checkpoint) against opensource models OpenLLaMA-1 and OpenLLaMA-2.
Task Metric HLAT OpenLLaMA-1 OpenLLaMA-2
MMLU_STEM accuracy (norm) 34.833 (3.895) 28.5 (3.735) 33.895 (3.869)
MMLU_humanities accuracy (norm) 43.632 (3.266) 30.909 (3.099) 42.328 (3.288)
MMLU_other (business, health, misc.) accuracy (norm) 43.691 (3.669) 31.738 (3.463) 44.766 (3.671)
MMLU_social sciences accuracy (norm) 45.773 (3.447) 31.858 (3.302) 46.179 (3.503)
bigbench_causal_judgement multiple_choice_grade 52.632 (3.632) 56.316 (3.608) 53.158 (3.63)
bigbench_date_understanding multiple_choice_grade 61.518 (2.536) 58.537 (2.568) 56.911 (2.581)
bigbench_disambiguation_qa multiple_choice_grade 35.271 (2.981) 29.845 (2.854) 36.434 (3.002)
bigbench_dyck_languages multiple_choice_grade 21.7 (1.304) 17.5 (1.202) 19.5 (1.254)
bigbench_formal_fallacies_syllogisms_negation multiple_choice_grade 50.592 (0.42) 49.761 (0.42) 49.965 (0.42)
bigbench_geometric_shapes multiple_choice_grade 15.32 (1.904) 24.791 (2.282) 18.663 (2.059)
bigbench_hyperbaton multiple_choice_grade 50.242 (0.224) 49.358 (0.224) 51.064 (0.224)
bigbench_logical_deduction_five_objects multiple_choice_grade 23.0 (1.884) 28.4 (2.019) 26.8 (1.983)
bigbench_logical_deduction_seven_objects multiple_choice_grade 17.714 (1.444) 15.714 (1.377) 19.0 (1.484)
bigbench_logical_deduction_three_objects multiple_choice_grade 36.333 (2.781) 36.0 (2.776) 39.333 (2.825)
bigbench_movie_recommendation multiple_choice_grade 63.2 (2.159) 46.8 (2.234) 44.6 (2.225)
bigbench_navigate multiple_choice_grade 51.9 (1.581) 48.8 (1.581) 50.3 (1.582)
bigbench_reasoning_about_colored_objects multiple_choice_grade 37.95 (1.085) 35.1 (1.068) 37.4 (1.082)
bigbench_ruin_names multiple_choice_grade 35.491 (2.263) 31.696 (2.201) 31.473 (2.197)
bigbench_salient_translation_error_detection multiple_choice_grade 18.737 (1.236) 22.044 (1.313) 16.934 (1.188)
bigbench_snarks multiple_choice_grade 46.409 (3.717) 49.724 (3.727) 43.094 (3.691)
bigbench_sports_understanding multiple_choice_grade 58.519 (1.57) 50.71 (1.593) 56.187 (1.581)
bigbench_temporal_sequences multiple_choice_grade 22.6 (1.323) 26.6 (1.398) 24.5 (1.361)
bigbench_tracking_shuffled_objects_five_objects multiple_choice_grade 18.64 (1.102) 19.52 (1.122) 18.16 (1.091)
bigbench_tracking_shuffled_objects_seven_objects multiple_choice_grade 13.771 (0.824) 13.029 (0.805) 12.743 (0.797)
bigbench_tracking_shuffled_objects_three_objects multiple_choice_grade 36.333 (2.781) 36.0 (2.776) 39.333 (2.825)
ARC Challenge accuracy 37.884 (1.418) 37.628 (1.416) 38.737 (1.424)
accuracy (norm) 40.188 (1.433) 38.567 (1.422) 41.382 (1.439)
ARC Easy accuracy 72.18 (0.92) 71.17 (0.929) 72.433 (0.917)
accuracy (norm) 66.709 (0.967) 68.855 (0.95) 69.739 (0.943)
HellaSwag accuracy 53.436 (0.498) 52.549 (0.498) 55.756 (0.496)
accuracy (norm) 71.251 (0.452) 69.189 (0.461) 74.517 (0.435)
OpenBookQA accuracy 28.8 (2.027) 29.8 (2.048) 29.8 (2.048)
accuracy (norm) 41.6 (2.206) 39.0 (2.183) 40.8 (2.2)
PIQA accuracy 76.496 (0.989) 75.68 (1.001) 78.727 (0.955)
accuracy (norm) 77.53 (0.974) 76.442 (0.99) 79.869 (0.936)
WinoGrande accuracy 68.114 (1.31) 66.693 (1.325) 65.904 (1.332)
NaturalQuestions exact match 22.632 (0.697) 22.327 (0.693) 20.139 (0.668)
TriviaQA exact match 55.06 (0.371) 55.556 (0.371) 53.912 (0.372)
BoolQ accuracy 72.508 (0.781) 70.459 (0.798) 72.416 (0.782)
GSM8K accuracy 9.401 (0.804) 5.08 (0.605) 5.231 (0.613)
Table 6. Evaluation results of HLAT on intermediate checkpoints. For Commonsense Reasoning (CR), we report the normalized accuracy. For Code, we report pass@100.
Tokens (B) MMLU BBH CR WK RC Math Code
120 28.49 (3.357) 32.824 (1.835) 50.761 (1.211) 21.405 (0.431) 61.468 (0.851) 1.365 (0.32) 15.85
200 24.759 (3.217) 33.919 (1.843) 53.14 (1.217) 23.572 (0.448) 64.098 (0.839) 2.426 (0.424) 18.90
400 26.066 (3.278) 33.948 (1.835) 55.208 (1.212) 27.89 (0.481) 56.575 (0.867) 2.274 (0.411) 20.12
600 27.936 (3.331) 33.904 (1.835) 55.489 (1.208) 28.434 (0.485) 67.187 (0.821) 2.35 (0.417) 21.95
800 27.947 (3.336) 34.701 (1.835) 56.17 (1.212) 30.183 (0.49) 68.257 (0.814) 2.881 (0.461) 23.78
1000 29.034 (3.391) 34.571 (1.842) 56.872 (1.213) 32.713 (0.504) 67.859 (0.817) 3.867 (0.531) 28.05
1200 33.367 (3.503) 34.991 (1.853) 57.458 (1.212) 33.476 (0.507) 69.174 (0.808) 4.17 (0.551) 27.44
1400 35.54 (3.546) 35.495 (1.849) 58.279 (1.208) 35.674 (0.521) 72.202 (0.784) 4.776 (0.587) 32.93
1500 37.701 (3.563) 35.823 (1.853) 58.862 (1.203) 35.878 (0.525) 71.346 (0.791) 5.762 (0.642) 32.31
1600 36.82 (3.532) 35.334 (1.835) 59.616 (1.204) 36.758 (0.526) 71.651 (0.788) 6.368 (0.673) 32.93
1800 41.318 (3.602) 36.565 (1.845) 59.455 (1.206) 38.846 (0.534) 72.508 (0.781) 9.401 (0.804) 34.15
Appendix B Evaluation Results on Intermediate Checkpoints
In Table 6 we list the evaluation results of HLAT on intermediate checkpoints for every 200 billions of training tokens. For Commonsense Reasoning, we report the normalized accuracy if exists. For Code, we report pass@100. We also plot the trends in Figure 5.
Figure 5. Performance progress of HLAT on intermediate checkpoints over individual evaluation tasks.
Appendix C Neuron Nemo Megatorn Package
We also experimented on AWS Neuron Nemo Megatron package [^10] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/libraries/nemo-megatron/index.html, which is adapted from Nemo (Harper et al., [n. d.]) for running on AWS Trainium  accelerators. In Figure 6, we plot both training loss and gradient norm of two runs with same configuration on NDTL and Neuron Nemo Megatron, respectively. Neuron Nemo Megatron provides same convergence as NDTL.
Figure 6. Convergence comparison of Neuron Nemo Megatron and NDTL. (a) The training loss vs number of tokens (in billions) seen by the model during training. (b) Gradient norm.

### v2

@@ V-E Truthfulness and Bias
We report the model’s truthfulness and bias using TruthfulQA [50] and CrowS-pairs [51]. TruthfulQA presents a collection of meticulously crafted questions spanning diverse domains such as health, law, finance, and even politics. These queries deliberately target areas where human intuition and personal biases can lead to incorrect responses, and measure an LLM’s resistance to misinformed or erroneous knowledge. CrowS-Pairs is a benchmark designed to probe LLMs for social biases across nine categories, including gender, religion, race/color, sexual orientation, age, nationality, disability, physical appearance and socioeconomic status. Each example is composed of a stereotype and an anti-stereotype.
TABLE V: Model Truthfulness and Bias evaluation. CrowS-pairs (CSP) uses percentage of stereotypes as metric and TruthfulQA (TQA) uses multiple choice accuracy as metric.
Dataset Size CSP ($\downarrow$) CSP TQA ($\uparrow$) TQA
Tasks - english french mc1 mc2
OpenLLaMA-1 7B 64.6 50.1 23.1 35.1
OpenLLaMA-2 7B 65.6 51.7 22.6 34.6
LLaMA-1 7B 53.7 47.5 22.0 34.1
LLaMA-2 7B 66.9 54.9 25.2 39.0
HLAT-7B 7B 65.2 54.5 23.6 37.2
LLaMA-1 65B 69.3 58.3 27.9 42.6
LLaMA-2 70B 69.8 63.5 30.6 44.8
HLAT-70B 70B 68.1 59.1 32.3 45.9
@@
We present the results in Table V with 0 shot inference. For TruthfulQA, we measure the multiple-choice score, and higher score shows better truthfulness. For CrowS-Pairs, it measures the percentage of models choosing answers of stereotypes, so lower scores indicates smaller bias. Overall, HLAT performs similar to other open-source models.
V-F Efficiency and Scalability
We describe the training efficiency in terms of Cost per 4-million tokens (CPT) and scalability reported in [52]. The CPT is defined as $\text{CPT}=\frac{C}{T\times 3600}\times 4\mathrm{e}^{6}$, where $C$ is instance cost per hour (\$21.50 for Trainium, and \$32.77 for GPU), $T$ is training throughput (tokens per second). CPT quantifies both the training speed and also hardware cost. We use this metric to compare training efficiency of Trainium and GPU.
@@ V-F Efficiency and Scalability
Figure 5: Normalized cost per 4 million tokens (CPT) for 7B and 70B models on AWS Trainium with various number of nodes. CPT of GPU baseline (p4d, 7B) with 4 nodes is normalized to 100%. 70B models ran into out-of-memory on 4 nodes.
For comparison, the GPU baseline is established using p4d.24xlarge instances and NeMo 23.08 [25] (available inside NeMo docker container with tag 23.08) software stack. Figure 5 plots the normalized CPT of training on Trainium and scaling. The Trainium CPT is normalized, such that the CPT of the GPU baseline (p4d, 7B) on 4 nodes is 100%. Overall, the training cost on trn1 is approximately 60% of GPU, and is consistent with the number of nodes. In addition, the CPTs on 70B models are roughly 10 times of those on 7B models.
V-G Model Limitation
@@
We note some limitations of HLAT in this section. Similar as other LLMs, HLAT suffers a set of limitations such as hallucinations, potential non-factual generations, biases, and toxicity [53]. For example, although comparable with other open-source pre-trained models, the bias of HLAT is still relative high on some subjects such as sexual orientation, physical appearance, religion, and socioeconomic (see Table V). This is partially due to the usage of publicly available datasets. More importantly, as a pre-trained model, HLAT has not gone through a supervised finetuning and human preference alignment. Those fine-tuning methods have been shown to be able to alleviate some limitations of pre-trained LLMs [9]. Another limitation is that our training is stopped after 1.8 trillion tokens. As is suggested by LLaMA-3 [10], HLAT may be able to further improve on certain tasks, such as math, world knowledge, MMLU, and coding, with more training tokens.
VI Best Practices & Future Directions
@@
Parallelism: NxDT supports TP up to 32 degrees and pipeline parallelism. For a 7B model, we found that the combination of TP=8 and PP=1 provides the highest training throughput, but not for HLAT-70B. So the optimal parallelism configuration varies with model sizes and architectures. To achieve the highest training throughput, parallelism configuration needs to be jointly optimized with choice of activation checkpointing method, gradient accumulation steps, and training precision, to balance memory and communication costs.
Training Precision: NxDT supports various training precision configurations, including full precision (FP32), BF16 with and without SR, standard mixed precision training, etc. Full precision training is often memory-wise infeasible for multi-billion LLMs. We compared multiple training strategies for HLAT: pure BF16, BF16 with SR and standard mixed precision training. Empirically, we found that training loss of pure BF16 diverges. BF16 with SR shows similar training loss as mixed precision on HLAT-7B model, but converges slower on HLAT-70B. We finally chose BF16 with SR for higher throughput on HLAT-7B, but standard mixed precision on HLAT-70B. For models of other sizes and architecture, warm-up study may be needed to decide the optimal training precision. Usually, the divergence can be observed in first few thousands of steps.
@@
Weight decay: We applied weight decay to all layers. Empirically, weight decay is not applied to normalization and bias layers [54]. In our experiment, we did not found much performance-wise difference of those two methods.
@@ VII Related Work
LLM pre-training: After the Transformer architecture [1] was introduced, BERT [54] was proposed to pre-train a language model on a large corpus of unlabeled data. Following the success of BERT model on various NLP tasks, many pre-trained language models are later introduced with different architectures and training methods, such as GPT-2 [55], RoBERTa [56], BART [57], and so on [6]. Studies later observed significant performance improvement of language models by increasing model size and training data [58]. Such abilities are further demonstrated in LLMs such as GPT-3 [7], PaLM [59], LLaMA [8, 9, 10], Falcon [60], Gemini [61], Phi [48], etc. Pre-trained on trillions of tokens, LLMs with tens or hundreds of billions parameters show remarkable ability in generating creative text contents, as well as a variety of downstream tasks, such as question answering, summarization, machine translation, programming, etc. [6].
@@
AI accelerators: Most models are trained on NVIDIA GPU accelerators, such as GPT [7, 62] and LLaMA [8, 9]. Falcon-180B [60] was trained on AWS SageMaker, with up to 4,096 A100 40GB GPUs using p4d instances. However, the landscape of hardware accelerators for deep learning training has blossomed in recent years, with established players like NVIDIA GPUs facing fierce competition from custom offerings like Google’s TPU and AWS Trainium. PaLM-2 [59] and OpenLLaMA [29] have demonstreated successful LLM pre-training on Google TPU. Recently, OLMo [49] is an open-source model developed by AI2. It has two models trained on AMD and Nvidia GPUs, separately. The two models have nearly identical performance on their evaluation suite by 2T tokens. AWS Trainium is a machine learning accelerator developed for deep learning training with high performance and cost-competitiveness. Our work is the first demonstration of end-to-end multi-billion LLM pre-trained on AWS Trainium. Ultimately, the optimal choice depends on the specific needs of the training task, with further research required to fully explore the potential of each accelerator and their possible convergence in future architectures.
VIII Conclusion
@@
In this paper, we pre-train HLAT, a family of 7 billion and 70 billion parameter large language models, using AWS Trainium over $\sim$1.8 trillion tokens. HLAT follows the decoder-only architecture and is trained with up to 256 Amazon EC2 trn1.32xlarge instances. We evaluate the performance of HLAT against popular open-source baseline models including LLaMA and OpenLLaMA on a variety of popular benchmarking tasks. We find that HLAT achieves model quality on par with these baseline models of similar sizes. This work demonstrates, for the first time, that AWS Trainium with NxDT is able to successfully pre-train high-quality LLMs with high efficiency and low cost.